OpenStudy (anonymous):

Find the exact value of sin-1(-0.5).

OpenStudy (anonymous):

pi/3 correct?

OpenStudy (anonymous):

no because \[\sin(\frac{\pi}{3})=\frac{1}{2}\] not \(-\frac{1}{2}\) but you have the right idea

OpenStudy (anonymous):

go the other way on the unit circle to \(-\frac{\pi}{3}\)

OpenStudy (anonymous):

ok so -pi/3

OpenStudy (anonymous):

it makes sense

OpenStudy (anonymous):

yup don't forget your answer has to be between \(-\frac{\pi}{2}\) and \(\frac{\pi}{2}\) which is why you are not allowed to say \(\frac{2\pi}{3}\) for example

OpenStudy (anonymous):

got :))

OpenStudy (anonymous):

good!